Solve yzdx-xzdy-(x^2+y^2)tan^-1(y/x)dz=0


1
Expert's answer
2021-06-07T19:33:17-0400

Question:-

"yzdx-xzdy-(x^2+y^2)tan^{-1}(\frac{y}{x})dz=0..........(1)"



Given

"yzdx-xzdy-(x^2+y^2)tan^{-1}(\frac{y}{x})dz=0..........(1)\\or\\
yzdx-xzdy=(x^2+y^2)tan^{-1}(\frac{y}{x})dz"


Lagrange's auxiliary equation

"\frac{dx}{yz}=\frac{dy}{-xz}=\frac{dz}{(x^2+y^2)tan^{-1}(\frac{y}{x})}"

"from \space I=II"

-xdx=ydy

y2+x2=c1............(2)


again


"yzdx-xzdy-(x^2+y^2)tan^{-1}(\frac{y}{x})dz=0\\
or
\\
ydx-xdy-(x^2+y^2)tan^{-1}(\frac{y}{x})\frac{dz}{z}=0"

let

"\\ \frac{y}{x}=t\\
diffrentiate\\
\frac{xdy-ydx}{x^2}=dt"

put the values and we get

"(-x^2dt)-(x^2+y^2)tan^{-1}(t)\frac{dz}{z}=0\\
-dt -(1+t^2) tan^{-1}(t) \frac{dz}{z}=0\\
or\\
\frac{dt}{(1+t^2)(tan^{-1}(t))}+\frac{dz}{z}=0\\
let 
\\tan^{-1}(t)=m\\
differentiate\\
\frac{dt}{(1+t^2)}=dm\\
put \space values,\\
\frac{dm}{m}+\frac{dz}{z}=0\\
integrate\\
log(mz)=log(c_2)\\
mz=c_2\\
again \space put \space values(m,t)\\
z .tan^{-1}(\frac{y}{x})=c_2........(3)
\\hence\\
answer\\
f(c_1,c_2)=0
\\f(x^2+y^2,z.tan^{-1}(\frac{y}{x})=0"